### v1

# Towards Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation: An Empirical Study

Abstract.
Multi-modal foundation models and generative AI have demonstrated promising capabilities in applications across various domains. Recently, Vision-language-action (VLA) models have attracted much attention regarding their potential to advance robotic manipulation. Despite the end-to-end perception-control loop offered by the VLA models, there is a lack of comprehensive understanding of the capabilities of such models and an automated testing platform to reveal their robustness and reliability across different robotic manipulation scenarios. To address these challenges, in this work, we present VLATest, a testing framework that automatically generates diverse robotic manipulation scenes to assess the performance of VLA models from various perspectives. Large-scale experiments are considered, including eight VLA models, four types of manipulation tasks and over 18,604 testing scenes. The experimental results show that existing VAL models still lack imperative robustness for practical applications. Specifically, the performance of VLA models can be significantly affected by several factors from the operation environments such as camera poses, lighting conditions and unseen objects. Our framework and the insights derived from the study are expected to pave the way for more advanced and reliable VLA-enabled robotic manipulation systems in practice.
Vision-Language-Action Models, Robotic Manipulation, Robustness, Empirical Study
$c$s: Software and its engineering Software defect analysisccs: General and reference Empirical studiesccs: Computer systems organization Robotics
1. Introduction
Robotic manipulation is widely regarded as one of the most important areas within cyber-physical systems (CPS). Over the past few decades, robotic manipulation and its applications have been implemented across various domains, such as industrial automation (Dzedzickis et al., 2021; Goel and Gupta, 2020), healthcare (Kyrarini et al., 2021; Holland et al., 2021), and logistics and warehousing (Shamout et al., 2022; Dhaliwal, 2020). With the rapid advancement of AI techniques, researchers and practitioners have been exploring the integration of AI for planning (Guo et al., 2023; Rayhan, 2023) and control (Soori et al., 2023; Wakchaure et al., 2023) in robotic manipulation. For instance, recent studies have employed deep reinforcement learning for robotics control to enhance the adaptability of the system against unseen scenarios (Han et al., 2023; Orr and Dutta, 2023).
Meanwhile, the emergence of foundation models, such as large language models (LLMs) and vision-language models (VLMs), has introduced new opportunities for advancing AI-enabled robotic manipulation (Singh et al., 2023; Ding et al., 2023). In particular, LLMs and VLMs have demonstrated promising capabilities to participate in the loop of robotics system development and operation to enhance corresponding reasoning, perception, and task-planning abilities  (Huang et al., 2023a; Liang et al., 2023b; Driess et al., 2023). In contrast to the general purpose-oriented LLMs and VLMs, vision-language-action (VLA) models were developed exclusively to generate robot actions for manipulation based on visual observations from cameras and task instructions provided in users’ natural language input. Techniques such as reinforcement learning from human feedback (RLHF) (Bai et al., 2022) further enhance these models’ ability to interpret human intent from natural language commands. As a result, it is now possible to let robotics perform manipulation tasks simply by prompting natural language commands into the VLA models. Additionally, the large-scale pre-training nature of such foundation models makes it possible to adapt to a diverse range of downstream manipulation tasks without task- and environment-specific niche design. A recent study demonstrates that the state-of-the-art (SOTA) VLA model, RT-2 (Brohan et al., 2023), can be extended to complex tasks such as reasoning, symbol understanding, and human recognition—tasks that were unseen during training (Brohan et al., 2023).
Figure 1. Architecture and workflow of a VLA model. (a) Generating robot actions at the first timestamp. (b) Generating robot actions at the k-th timestamp.
However, the data-driven nature of VLA models makes them difficult to interpret, which hinders the safety and applicability of VLA models to more robotics applications in practice. To enhance the reliability and trustworthiness of VLA models for robotic manipulation, the development of quality assurance techniques such as testing, debugging, and repairing has become an urgent need in both industrial and academic communities. In particular, comprehensive and adequate testing is essential for throughout assessing a VLA model’s capabilities and limitations. Unfortunately, current VLA models are typically evaluated using only a limited set of hand-crafted scenes, where the comprehensiveness and effectiveness of such scenes are often inadequate, which subsequently leaves the general performance and behavior characteristics of VLA models largely unexplored. Moreover, unlike conventional LLMs, which only process Natural Language (NL) text inputs, VLA models are empowered by the multimodality perception ability; therefore, when deploying VLA models in real-world scenarios, visual factors such as obstacles, lighting conditions, and camera angles may significantly impact their performance. Despite this, there is currently a lack of deep understanding regarding the robustness of VLA models against the variations among such environmental factors.
To address this gap, we propose VLATest, one of the first testing frameworks specifically designed for evaluating VLA models in robotic manipulation. We introduce a set of ten testing operators and design a generation-based fuzzing framework that automatically produces testing scenes to assess the performance and robustness of VLA models in specific robotic manipulation tasks. VLATest is implemented within the Maniskill2 simulation environment (Gu et al., 2023). By leveraging VLATest, we facilitate efficient testing and enable comprehensive evaluation of VLA models under varying conditions. Based on VLATest, we performed a large-scale empirical study to evaluate the overall performance and robustness of popular VLA models. A set of research questions are proposed to examine key factors in reliable robotic manipulation with VLA: basic performance, task complexity, perception robustness, and OOD (out-of-distribution) robustness. Specifically, we investigated the following five research questions:
  • RQ1
    How do VLA models perform in popular robotic manipulation tasks?
  • RQ2
    How does the number of obstacles affect a VLA model’s performance?
  • RQ3
    Does the change in lighting conditions affect a VLA model’s performance?
  • RQ4
    Does the change of camera pose affect a VLA model’s performance?
  • RQ5
    How robust do VLA models perform against unseen objects?
To investigate these research questions, we generated 18,604 testing scenes across four different robotic manipulation tasks. We selected seven popular publicly available VLA models: RT-1-1k, RT-58k, RT-400k (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo-small, Octo-base (Team et al., 2024), and OpenVLA-7b (Kim et al., 2024). The experiments included a total of 78,604 rounds of simulation execution, which took more than 400 GPU hours.
Our analysis shows that current VLA models exhibit subpar performance across the four robotic manipulation tasks studied. As the number of obstacles increases, it becomes more challenging for these models to accurately locate and manipulate the correct objects. Additionally, we observed that the subject VLA models lack robustness when subjected to changes in lighting conditions and camera angles, performing much worse compared to the default settings. Meanwhile, we found that VLA models with large-scale pre-training generally show better robustness against such changes. Finally, we found that these VLA models struggle significantly with unseen objects, exhibiting a significant drop in performance compared to manipulating seen objects. Our findings provide practical guidelines for developers working with VLA models on specific robotic manipulation tasks and underscore the need for more advanced VLA models, as well as novel quality assurance techniques to enhance the robustness and reliability of the VLA models.
In summary, this paper makes the following contributions:
  • A testing framework. We designed and implemented a generation-based fuzzing framework, VLATest, to test VLA models by incorporating various operators in robotic manipulation testing scenes.
  • An empirical study. Based on VLATest, we conducted a large-scale empirical study to evaluate the performance and robustness of seven popular VLA models across four robotic manipulation tasks under various conditions, including variations in the number of obstacles, lighting conditions, camera poses, and the manipulation of both seen and unseen objects.
  • Implications. We discussed the challenges and limitations of current VLA models, along with the implications and future opportunities for enhancing their robustness and reliability.
2. Background: Vision-Language-Action Models
Vision-Language-Action (VLA) models (Brohan et al., 2022; Padalkar et al., 2023; Team et al., 2024; Kim et al., 2024) are a type of deep neural network that takes natural language input from the user as task instructions and visual input from a camera as observations. The output of a VLA model is a set of actions to achieve the designated manipulation according to the task instructions, such as moving a robotic arm’s joints and opening the gripper.
Figure 2. Operators used for testing VLA models in VLATest.
2.1. VLA Model for Robotic Manipulation
Fig. 1 illustrates the common architecture of a VLA model. Given a natural language input $T$ and an image input $I_{1}$ at the first timestamp, the VLA model’s natural language tokenizer and visual encoder first project $T$ and $I_{1}$ into sets of tokens, $\mathbf{T}=\{t_{1},\dots,t_{m}\}$ and $\mathbf{I}_{1}=\{i_{11},\dots,i_{1n}\}$, respectively. These tokens are then concatenated and fed into a transformer model to predict action token(s) $\mathbf{A}_{1}$. An action head layer then de-tokenizes $\mathbf{A}_{1}$ into robot action values $\Delta_{x1},\Delta_{\theta 1},\Delta_{grip1}]$, where $\Delta_{x1}\in\mathbb{R}^{3}$, $\Delta_{\theta 1}\in\mathbb{R}^{3}$, and $\Delta_{grip1}\in\mathbb{R}^{1}$ denote the translation ($x,y,z$ movement), rotation, and gripper actions of the end-effector that should be performed after the observation, respectively.
After the robot performs these actions, the action tokens $\mathbf{A}_{1}$ and a new set of image tokens $\mathbf{I}_{2}=\{i_{21},\dots,i_{2n}\}$, based on the visual observation at the next timestamp, are concatenated with $\mathbf{T}$ and $\mathbf{I}_{1}$. A new input sequence $\{\mathbf{T},\mathbf{I}_{1},\mathbf{A}_{1},\mathbf{I}_{2}\}$ is then fed into the transformer to predict new action tokens $\mathbf{A}_{2}$. The VLA model continues this process until the task is completed or the pre-defined maximum number of steps is exceeded.
2.2. Training and Evaluation of the VLA Model
Training. There are typically two approaches to training a VLA model: (1) training from scratch and (2) fine-tuning a general-purpose vision-language model (VLM). Training from scratch involves building a VLA model directly on robot demonstration data. This approach is commonly applied for models with relatively small-scaled architectures and limited computational resources, such as RT-1 (Brohan et al., 2022) and Octo (Team et al., 2024), which typically have fewer than 100 million parameters. By contrast, fine-tuning from an existing VLM leverages the flexibility of re-using a larger model (more than 1 billion parameters), such as Llava (Liu et al., 2024a; Kim et al., 2024), which has been pre-trained on massive amounts of image and text data from diverse domains. The large-scale pre-training of these VLMs may enhance the generalizability of the fine-tuned VLA models, particularly when manipulating unseen objects and tasks.
Evaluation. A VLA model is usually evaluated by measuring its performance on specific skills, such as picking up an object in a scene. To evaluate a particular skill, developers must first create an adequate testing scene, which involves configuring target objects, obstacles, and environmental factors like lighting conditions. Additionally, a text prompt should be meticulously crafted before executing the VLA model to perform the manipulation task. To assess the model’s performance, developers usually design a set of evaluation metrics. For instance, when performing the task “picking up an object”, the metrics include: (1) whether the robot grasps the correct object, (2) whether the object is successfully lifted, and (3) whether the robot can sustain lifting the object for a short period. The execution of the VLA model and the robot can occur in either a simulated or real-world environment. In real-world scenarios, manual labeling is required to compute these metrics, whereas, in simulated environments, the evaluation process can be automated.
3. VLATest
In this section, we first introduce the operators included in the testing scene for a robotic manipulation task. Next, we describe the algorithm used to generate a testing scene in VLATest.
3.1. Operators
We consider four categories of testing operators in VLATest as shown in Fig. 2, resulting in a total of ten testing operators.
  • Target object(s). A target object refers to an object that is required to be manipulated in a task (e.g., an object to pick up) or an object that serves as a reference in a task (e.g., an object to be placed on). For each target object, we consider three different operators: (1) type of object (e.g., Pepsi can), (2) position, and (3) orientation.
  • Obstacles. Different from target object(s), obstacles are typically not part of the task, and the robot is expected to avoid contacting these objects during manipulation. In addition to the three operators for target object(s), we also consider the number of obstacles as a distinct operator in the obstacles category.
  • Lighting. The lighting condition affects the rendering of images captured by the camera, thereby influencing the visual input of the VLA models. In VLATest, we consider lighting intensity as one of the operators.
  • Camera. The camera’s pose can also affect the visual input of the VLA models. We consider its position and orientation as two separate operators in VLATest.
3.2. Testing Scene Generation
We summarize the algorithm for generating a testing scene in Algorithm 1. Intuitively, there are four distinct steps: (1) sampling semantically valid target object(s), (2) sampling obstacles (optional), (3) mutating lighting conditions (optional), and (4) mutating camera poses (optional).
Sampling semantically valid target object(s). For a robotic manipulation task, we first randomly sample target object(s) without replacement from the object database $\mathcal{O}$ (Lines 5-9). We also check if the selected objects are semantically valid according to the task requirements (Line 4). For example, in a task requiring object A to be placed over object B, if object B is a ball, object A can never be stably placed on it. Such cases are considered semantically invalid. After selecting the target object(s), we proceed to sample their positions and orientations (Line 7). In our empirical study, we used a random sampler for pose selection. If there are multiple target objects, we ensure they remain far enough apart (>safe_dist>safe\_dist) to avoid generating invalid testing scenes. For instance, in the previous example, if object A is already placed on object B during the sampling of their positions, the testing scene becomes invalid.
Input: an object database $\mathcal{O}$, the robotic manipulation task $T$, the number of target objects $N_{target}$, the number of obstacles $N_{obstacles}$, the lighting mutation option lighting_flaglighting\_flag, and the camera mutation option camera_flagcamera\_flag, the minimum distance between objects safe_distsafe\_dist
Output: a testing scene
1 $\mathcal{O}~\leftarrow~\mathsf{random\_shuffle}(\mathcal{O}$);
2 $O_{target},O_{obstacles}~\leftarrow~\emptyset,\emptyset$;
[…]
4 while not $\mathsf{semantic\_valid}(O_{target},T)$ do
[…]
6     $obj~\leftarrow~\mathcal{O}\mathsf{.pop()}$;
7     $pos,ori~\leftarrow~\mathsf{pose\_sampler}(O_{target},safe\_dist)$;
8     $O_{target}~\leftarrow~O_{target}\bigcup(obj,pos,ori)$;
9
10  end for
11
12 end while
13if $N_{obstacles}>0$ then
[…]
15     $obj~\leftarrow~\mathcal{O}\mathsf{.pop()}$;
16     $pos,ori~\leftarrow~\mathsf{pose\_sampler}(O_{target}\bigcup O_{obstacles},$ safe_dist)safe\_dist);
17     $O_{obstacles}~\leftarrow~O_{target}\bigcup(obj,pos,ori)$;
18
19  end for
20
21 end if
22if lighting_flaglighting\_flag then
23  $intensity~\leftarrow~\mathsf{mutate\_lighting}(intensity)$;
24
25 end if
26if camera_flagcamera\_flag then
27  $cam_{pos},cam_{ori}~\leftarrow~\mathsf{mutate\_camera}(cam_{pos},$ $cam_{ori})$;
28
29 end if
30return $\{O_{target},O_{obstacles},intensity,(cam_{pos},cam_{ori})\}$;
Algorithm 1 Testing scene generation in VLATest.
Sampling obstacles. If the number of obstacles $N_{obstacles}$ is greater than 0 (Line 11), we further randomly sample obstacle objects from $\mathcal{O}$ without replacement (Lines 12-16). For each obstacle, we follow the same procedure as used for sampling the target objects to assign their position and orientation (Line 14). In our empirical study, we utilized a random sampler to generate the position and orientation for each obstacle.
Mutating lighting conditions and camera poses. For lighting conditions, we generate a factor $\alpha$ to mutate the lighting intensity value (Lines 18-20). For the camera, we adjust its position by a distance $d$ and rotate it by an angle $\theta$ (Lines 21-23). Note that both $d$ and $\theta$ are kept small to ensure that the entire scene remains within the camera’s view. In our empirical study, we used a random generator to generate $\alpha$, $d$, and $\theta$ within specified ranges.
4. Empirical Study
Based on VLATest, we conducted an empirical study to investigate the performance of SOTA VLA models. In this section, we first list our research questions, followed by the empirical setup, which includes subject VLA models, subject robotic manipulation tasks, prompt template, and implementation details.
4.1. Research Questions
Our empirical study investigates the following research questions to examine the key factors in reliable robotic manipulation: basic performance (RQ1), task complexity (RQ2), perception robustness (RQ3 & RQ 4), and OOD robustness (RQ5).
  • RQ1. How do VLA models perform in popular robotic manipulation tasks? This research question seeks to evaluate the performance of SOTA VLA models in various robotic manipulation tasks. While previous studies (Brohan et al., 2022; Team et al., 2024) typically rely on a few hand-crafted test cases, we use VLATest to generate a large number of test cases. This approach enables a more comprehensive assessment of SOTA VLA models’ performance, providing insights into current challenges and opportunities.
  • RQ2. How does the number of obstacles affect a VLA model’s performance? Intuitively, the presence of more obstacles (i.e., objects unrelated to the assigned manipulation task) increases the complexity of a robotic manipulation task for VLA models. However, it remains unclear whether there is an upper limit to the task complexity that a VLA model can handle effectively. To address this, we conduct controlled experiments to help practitioners understand how VLA models perform as the number of obstacles varies.
  • RQ3. Does the change in lighting conditions affect a VLA model’s performance? When deploying robotics and VLA models in real-world environments, external conditions can vary significantly. Ideally, a VLA model should be robust against different lighting conditions, such as varying illumination intensities. This research aims to offer practitioners practical guidelines on the robustness of VLA models under diverse lighting setups.
  • RQ4. Does the change of camera pose affect a VLA model’s performance? Since VLA models are pre-trained on large-scale vision datasets, they are expected to demonstrate robustness when input images are captured from various angles. This research question seeks to determine whether and to what extent current VLA models maintain robustness when operating with different camera poses.
  • RQ5. How robust do VLA models perform against unseen objects? Objects that were unseen in the robotic demonstration data may present when deploying robotics and VLA models in practical scenarios. It is unclear whether the large-scale pre-training of VLA models can make it generalizable to these unseen objects. Also, there is currently limited understanding of the performance gap between seen-object and unseen-object robotic manipulation tasks. To address this, we investigate this research question by leveraging an external object database to assess the limitations and challenges.
4.2. Subject VLA Models
To investigate our research questions, we studied four different series of open-source VLA models: RT-1 (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo (Team et al., 2024), and OpenVLA (Kim et al., 2024). For RT-1 (Brohan et al., 2022), we studied three publicly available variants: RT-1-1k, RT-1-58k, and RT-1-400k  [^1] The number denotes the training steps taken for training an RT-1 model.. We also studied two variants for Octo (Team et al., 2024): Octo-small and Octo-base. Note that, we were unable to study the SOTA VLA model, RT-2 (Brohan et al., 2023), as it has not yet been made publicly available. Thus, we included a total of seven subject VLA models in our empirical study. Table 1 summarizes the characteristics of these models, and we provide more details for each model below.
Table 1. Overview of subject VLA models.
VLA Models RT-1 RT-1-X Octo-small Octo-base OpenVLA-7b
1k 58k 400k
Model Size 35M 35M 27M 93M 7.6B
Release Date Dec. 2022 Oct. 2023 Dec. 2023 Dec. 2023 Jun. 2024
  • RT-1 (Brohan et al., 2022) model was released by Google Research. The RT-1 model includes a FiLM EfficientNet-B3 (Perez et al., 2018; Tan and Le, 2019) pre-trained on the ImageNet dataset (Russakovsky et al., 2015) for tokenizing visual inputs and language instructions before connecting to a transformer to generate robotic manipulation actions. The RT-1 model was trained on an unpublished set of 130k robot demonstrations collected by Google.
  • RT-1-X (Padalkar et al., 2023) model was released by Google DeepMind. RT-1-X shares the same model architecture as RT-1. However, RT-1-X was trained on the open-source dataset, Open X-Embodiment (Padalkar et al., 2023), which includes 160k robot demonstrations collected from $2$ different robots.
  • Octo (Team et al., 2024) model, released by UC Berkeley, includes a ViT model (Dosovitskiy et al., 2020) as its backbone transformer. It introduces the “readout” token, allowing developers to flexibly add new observation inputs or action output heads to the model during downstream fine-tuning. Octo was trained on a subset of the Open X-Embodiment dataset, which includes about 65k robot demonstrations. Octo-small (27M parameters) and Octo-base (93M parameters) are two variants of Octo, utilizing ViT-S and ViT-B as their backbone transformers, respectively.
  • OpenVLA-7b (Kim et al., 2024) is the most recent VLA model released by Stanford University. OpenVLA-7b includes a 600M-parameter visual encoder consisting of pre-trained SigLIP (Zhai et al., 2023) and DinoV2 (Oquab et al., 2023) models and a 7B-parameter Llama 2 (Touvron et al., 2023) as the language model backbone. OpenVLA-7b was then fine-tuned on the Open X-Embodiment dataset.
4.3. Robotic Manipulation Tasks
We included four different robotic manipulation tasks in our study. Now we briefly introduce each of them. We also refer to the supplementary materials for the video demo of each task.
  • Task 1: Pick up an object. This task requires a VLA model to identify the target object and generate the appropriate control signals to grasp and lift it. To succeed, the robotic must grasp the correct object and lift it at least 0.02 meters for five consecutive frames.
  • Task 2: Move object A to object B. This task requires a VLA model to first identify the source object (A) and then output the corresponding control signals to move it near the target object (B). To succeed, the robotic must move the correct object to within 0.05 meters of the target object.
  • Task 3: Put object A on top of object B. Different from Task 2, this task requires the VLA model to output control signals that could stack object A on top of object B. To succeed, object A should be placed stably on top of object B.
  • Task 4: Put object A into object B. Different from Task 3, this task requires the VLA model to generate control signals that place object A inside object B (e.g., into a kitchen sink or a basket). To succeed, object A must be completely inside object B.
4.4. Prompt Templates
For each task, we follow the previous work (Brohan et al., 2022; Padalkar et al., 2023; Kim et al., 2024; Li et al., 2024b) to use the standard prompt template for each task: (1) pick up [object name], (2) move [object name] near [object name], (3) put [object name] on [object name], and (4) put [object name] into [object name].
4.5. Implementation Details
We conducted all experiments on a server with an AMD 5955WX CPU and two NVIDIA RTX A6000 GPUs. The operating system is 64-bit Ubuntu 20.04 LTS with Python 3.10 and CUDA 12.2. We implemented VLATest with Maniskill2 simulation environments (Gu et al., 2023; Li et al., 2024b). We used two object databases: (1) the default object database in Maniskill2 ($N=18$) in RQ1 ~ RQ4, and (2) YCB object database ($N=56$) in RQ5. The average execution time of one manipulation task was around 19.8 seconds. Our empirical study took about 410 GPU hours in total.
5. Results
5.1. RQ1: How Do VLA Models Perform in Popular Robotic Manipulation Tasks?
To investigate this research question, we leveraged VLATest to generate 1,000 scenes for each of the four robotic manipulation tasks by randomly selecting target object(s) and sampling 0 to 3 obstacles. We also randomly assigned poses to these objects. To avoid collision overlaps, we maintained a minimum distance of 0.15 meters between objects during the assignment. The default lighting setups and camera poses were used for this RQ.
Table 2. (RQ1) Performance of seven subject VLA models on different manipulation tasks. The top-1, top-2 and top-3 success rates for each step and task are highlighted, respectively.
VLA Models Task 1: Pick Up Task 2: Move Near Task 3: Put On Task 4: Put In
Grasp Lift Success Grasp Move Success Grasp Move Success Grasp Move Success
RT-1-1k 10.1% 1.2% 0.7% 9.6% 3.2% 1.4% 0.4% 0.0% 0.0% 1.1% 0.0% 0.0%
RT-1-58k 44.5% 32.6% 28.2% 38.6% 25.2% 10.9% 0.3% 0.0% 0.0% 1.3% 0.0% 0.0%
RT-1-400k 48.4% 41.0% 34.4% 38.8% 23.7% 9.4% 10.4% 1.3% 0.5% 8.9% 0.1% 0.1%
RT-1-X 34.0% 25.8% 19.5% 25.4% 15.2% 5.8% 17.2% 4.4% 2.3% 17.8% 0.7% 0.4%
Octo-small 9.0% 2.0% 0.8% 8.9% 4.1% 1.5% 27.5% 4.6% 2.2% 34.6% 1.1% 1.1%
Octo-base 2.3% 0.4% 0.0% 4.9% 1.5% 0.6% 19.1% 2.6% 1.2% 32.5% 1.1% 1.1%
OpenVLA-7b 15.1% 7.1% 5.9% 43.0% 23.3% 12.7% 36.8% 5.4% 2.1% 19.9% 1.1% 1.1%
Avg. 23.3% 15.7% 12.8% 25.0% 13.7% 6.0% 16.0% 2.6% 1.2% 16.6% 0.6% 0.5%
Figure 3. (RQ2) VLA performance vs. the number of obstacles.
Table 2 presents the performance of seven VLA models on these tasks and scenes. Overall, we find that current VLA models do not perform well in the four studied robotic manipulation tasks. The average success rates across the seven VLA models for the four tasks are 12.4%, 6.0%, 1.2%, and 0.5%, respectively. In Task 1, the best-performing VLA model, RT-1-400k, succeeded in only 34.4% of the test scenes. We also observed significant performance drops in Task 2, Task 3, and Task 4 compared to Task 1. The best-performing VLA model in Task 2, OpenVLA-7b, completed only 12.7% of the test scenes. In Task 3 and Task 4, the best-performing VLA models (RT-1-X and Octo-small) achieved success rates of just 2.2% and 2.1%, respectively. One possible reason is that these three tasks are more complex than Task 1, as they require multiple steps of reasoning (i.e., identifying the source and target objects and their positions before generating the corresponding control signals). Among the seven VLA models studied, we did not find any model that performed significantly better than the others across different tasks.
All VLA models exhibit subpar performance in the four studied robotic manipulation tasks, particularly in those that require identifying multiple target objects (i.e., Task 2, Task 3, and Task 4). These results suggest that the development of VLA models is still in its early stages, as they are far from being ready for deployment in real-world scenarios.
To understand the rationale behind the subpar performance of these VLA models, we broke down each testing scene according to the different steps required to complete a task successfully. Specifically, we measured the success rates of each step in a testing scene for the four tasks, as presented in Table 2. For instance, Task 1 requires the robot to (1) grasp the target object (the column Grasp in Table 2), (2) lift the object significantly (the column Lift in Table 2), and (3) continue lifting for five consecutive frames (the column Success in Table 2). In all four tasks, we found that the success rates dropped significantly between steps. For example, in Task 1, the average success rate of correctly grasping the target object is 23.3%. The success rates for lifting and maintaining the lift then drop to 15.7% and 12.4%, respectively. These results indicate that the current VLA models struggle to interpret natural language instructions that require the robot to perform multiple sequential actions.
In Task 2, Task 3, and Task 4, VLA models successfully picked up source objects in 16% to 25% of scenes. However, they then failed to identify the target objects, with only 0.6% to 13.7% of objects being correctly moved and 0.5% to 6.0% of objects being correctly placed. To improve, one could consider the idea of chain-of-thought prompting in LLMs. Specifically, the complex task instructions could be decoupled into individual action steps, allowing the VLA model to be prompted step-by-step.
Current VLA models cannot successfully execute designated task instructions that require multiple steps of action, highlighting the urgent need to improve their capabilities in accurately interpreting natural language prompts.
5.2. RQ2: How Does the Number of Obstacles Affect a VLA Model’s Performance?
To investigate RQ2, we used VLATest to generate 100 scenes for each task with a fixed number of $n$ obstacles, where $n$ was set to 0, 1, 2, 3, and 4, resulting in a total of 500 scenes per task. We used the default lighting conditions and camera poses when investigating this research question.
Figure 4. (RQ2) The number of successful testing scenes at different steps vs. the number of obstacles.
The results are depicted in Fig. 3. In the first two tasks (i.e., Task 1 and Task 2), we observed that the VLA model’s success rate decreased as the number of obstacles increased. The average success rates across different VLA models and scenes dropped from 17.3% to 8.3% and from 8.3% to 1.1% when increasing the number of obstacles from 0 to 4 in Task 1 and Task 2, respectively. However, in Task 3 and Task 4, we did not find a similar pattern. This may be because all VLA models performed poorly even without any obstacles ($n=0$). The average success rates across different VLA models and scenes were 1.2% and 0.7% when $n=0$ and 1.1% and 0.4% when $n=4$ for Task 3 and Task 4, respectively.
The number of obstacles affects the VLA model’s performance, indicating that VLA models become unreliable in more complex environments. When there were four obstacles presented in the scene, the VLA models only passed 8.2%, 2.3%, 1.0%, and 0.4% of the test scenes in Task 1, Task 2, Task 3, and Task 4, respectively.
Similar to RQ1, we also analyzed each testing scene and its success at each individual step. We counted the number of testing scenes that succeeded at each step and presented the results in Fig. 4. We found that the major challenge faced by the current VLA models lies in their inability to identify the correct object to manipulate.
In all four tasks, we found that the success rates for grasping the target object decreased as the number of obstacles increased. These results indicate that when there are multiple obstacles, the VLA models struggle to locate the correct object. Specifically, without any obstacles, the VLA models successfully located the object to manipulate in 199, 196, 137, and 152 out of 700 cases for Task 1, Task 2, Task 3, and Task 4, respectively. However, with four obstacles, the VLA models successfully located the correct object in only 107, 152, 78, and 65 out of 700 cases for the four tasks, respectively. The average success rates of grasping across VLA models dropped to 53.7%, 77.6%, 47.4%, and 42.8% for four tasks, respectively.
With multiple obstacles in the scene, the VLA models’ ability to accurately locate the correct object to manipulate can be significantly affected.
Fig. 5 shows two different testing scenes of Task 1. In both scenes, the VLA model, RT-1-X, was asked to pick up the Pepsi can. When there was no obstacle (Fig. 5(a)), RT-1-X was able to successfully grasp and lift the coke can. However, when there were two other obstacles (i.e., 7up can and Redbull can), RT-1-X failed to locate the Pepsi can and eventually picked up one of the obstacles.
(a) No obstacles
(b) Two obstacles
Figure 5. RT-1-X perform Task 1 with different number of obstacles.
5.3. RQ3: Does the Change in Lighting Conditions Affect a VLA Model’s Performance?
To answer this research question, we first collected all the successfully executed scenes for each VLA model in RQ1, resulting in a total of 1,434 passed test cases. We then re-executed these test cases three times after randomly increasing or decreasing the lighting intensities ($N=4,302$). Specifically, we randomly sampled a factor $\alpha\in(1,20]$ for increasing or $\alpha\in[1/20,1)$ for decreasing the lighting intensity. This factor was then multiplied by the default lighting intensity used in RQ1. We manually checked the cases when $\alpha$ was set to $20$ and $1/20$ and confirmed that the images captured under such lighting conditions were still recognizable by the humans.
Table 3. (RQ3) Performance of subject VLA models under default (Def.) and mutated (Mut.) lighting conditions. Each cell represents the number of successfully passed test scenes by different VLA models in different tasks. The top-1, top-2 and top-3 robust VLA models are highlighted, respectively.
VLA Models Task 1 Task 2 Task 3 Task 4 Overall
Pick Up Move Near Put On Put In
Def. Mut. Def. Mut. Def. Mut. Def. Mut. Def. Mut. Mut./Def. (%)
RT-1-1k 7 0.7 14 14.0 0 0 21 14.7 70.0%
RT-1-58k 282 142.0 109 109.0 0 0 491 251.0 51.1%
RT-1-400k 344 163.0 94 92.0 5 5.0 1 1.0 444 261.0 58.8%
RT-1-X 195 77.3 58 57.0 22 20.7 4 3.0 279 158.0 56.6%
Octo-small 8 0.0 15 2.3 21 6.0 10 1.0 54 9.3 17.2%
Octo-base 0 6 1.7 11 3.0 11 0.7 28 5.4 19.3%
OpenVLA-7b 59 21.0 127 127.0 20 20.0 11 11.0 217 169.0 77.9%
Tot. 895 404.0 423 403.0 79 54.7 37 16.7 1434 878.4 61.3%
  • $\dagger$ Averaged results over three mutations.
We present the experiment results in Table 3. Overall, we found that randomly perturbing the lighting conditions significantly affected the performance of the VLA models. Out of 1,434 passed test cases with default lighting conditions, only 878.4 test cases (average results over three mutations for each test case) could still be successfully executed with perturbed lighting conditions. The perturbed lighting conditions had a significant impact on Task 1, Task 3, and Task 4, where only about half of the test cases could still be passed after the perturbation. In Task 2, however, five out of seven VLA models exhibited robust performance despite changes in lighting conditions.
Among the seven VLA models, we found that OpenVLA-7b was the most robust, with 77.9% of the test cases still being passed under mutated lighting conditions. A plausible explanation is that OpenVLA-7b leverages two vision models (i.e., SigLIP (Zhai et al., 2023) and DinoV2 (Oquab et al., 2023)) pre-trained on a diverse set of images and fine-tuned on the LLaVA 1.5 data mixture (1 million images and texts). As a result, OpenVLA-7b can effectively interpret images captured under varying lighting conditions. In contrast, Octo-small and Octo-base were trained solely on images captured from robot demonstrations, making them less robust when faced with lighting conditions different from the default ones. Only 17.2% and 19.3% of the test cases could still be successfully executed after changing the lighting conditions, respectively.
Overall, the VLA models are not sufficiently robust against changes in lighting conditions. Only 61.3% of the test cases could still be successfully executed under mutated lighting conditions. Among the seven VLA models, OpenVLA-7b demonstrated the highest robustness to changes in lighting conditions.
We further investigated the extent to which the VLA models’ performance could be affected by increasing or decreasing lighting intensities. Fig. 6(a) and Fig. 6(b) present the results.
(a) Increasing the lighting intensity
(b) Decreasing the lighting intensity
Figure 6. VLA performance vs. different lighting intensities.
Increasing the lighting intensity ($\alpha>1$). Even when the lighting intensity is increased by less than 2.5 times (i.e., $\alpha<2.5$) of the default value, the success rate of the VLA models immediately drops to around 0.7. These results indicate that even a small increase in lighting intensity can significantly degrade the performance of the VLA models. As the factor $\alpha$ increases, the success rate of the VLA models continues to decrease. In extreme cases (e.g., $\alpha>7.5$), the VLA models succeed in less than half of the scenes.
The performance of the VLA models degrades with increasing lighting intensity. When the lighting intensity exceeds 8 times of the default value, the VLA models succeed in only about 40% of the testing scenes that could be passed under default lighting conditions.
Decreasing the lighting intensity ($\alpha<1$). Similar to increasing the lighting intensity, the VLA models’ performance is also affected by even a small perturbation (e.g., $\alpha>0.9$) when decreasing the intensity. The VLA models’ performance degrades with decreasing $\alpha$. However, compared to increasing the lighting intensity, decreasing it has less effects on the models’ performance. When $\alpha<0.2$, the VLA models can still pass around 60% of the testing scenes.
The effect of decreasing the lighting intensity is less significant than increasing it on the VLA models’ performance. Even when the lighting intensity is reduced to 0.15 of its default value, the VLA models can still pass 60% of the test cases.
5.4. RQ4: Does the Change of Camera Pose Affect a VLA Model’s Performance?
Similar to RQ3, we re-executed the 1,434 passed test cases collected from RQ1 three times after randomly moving and rotating the camera from its default position ($N=4,302$). Specifically, we rotated the camera around each axis by an angle randomly sampled between $5^{\circ}$ and $5^{\circ}$. Additionally, we randomly moved the camera away from its center by a distance randomly sampled between 0 and 5 cm. We manually checked the corner cases where the camera was rotated by $5^{\circ}$ and moved by 5 cm and confirmed that the images captured with these camera angles still included the entire scene and objects. We report our experiment results in Table 4
Table 4. (RQ4) Performance of subject VLA models with default (Def.) and mutated (Mut.) camera poses. Each cell represents the number of successfully passed test scenes by different VLA models in different tasks. The top-1, top-2 and top-3 robust VLA models are highlighted, respectively.
VLA Models Task 1 Task 2 Task 3 Task 4 Overall
Pick Up Move Near Put On Put In
Def. Mut. Def. Mut. Def. Mut. Def. Mut. Def. Mut. Mut./Def. (%)
RT-1-1k 7 0.7 14 4.3 0 0 21 5.0 23.8%
RT-1-58k 282 121.3 109 27.0 0 0 491 39.3 30.2%
RT-1-400k 344 180.3 94 21.3 5 0.0 1 0.7 444 202.3 45.6%
RT-1-X 195 43.0 58 13.3 22 0.7 4 0.0 279 57.0 20.4%
Octo-small 8 0.3 15 3.3 21 1.0 10 0.3 54 4.9 9.1%
Octo-base 0 6 0.7 11 1.7 11 0.3 28 2.7 9.6%
OpenVLA-7b 59 20.7 127 44 20 3.0 11 0.3 217 68.0 31.3%
Tot. 895 366.3 423 113.9 79 6.4 37 1.6 1434 488.2 34.0%
  • $\dagger$ Averaged results over three mutations.
Overall, we found that the VLA model’s performance was greatly affected by the mutated camera poses. With mutated camera poses, the subject VLA models only passed 34.0% of the test cases that could be passed with default camera poses. These results indicate that the current VLA models are very sensitive to the camera’s extrinsic calibration results. When deploying VLA models, to achieve compatible performance, the developers need to carefully set up the camera poses, as in the robot demonstration data used for training. However, these may limit the generalizability of the VLA models. Future work may consider data augmentation to improve the VLA model’s robustness under different camera settings to enhance the model’s generalizability.
Current VLA models are not robust against mutated camera poses, resulting in degraded performance when the visual input is captured from an angle different from the default one. When rotating the camera for a maximum angle of 5 and moving it for a maximum distance of 5cm, the performance dropped to 34.0% of the performance on average with default camera poses.
Among seven subject VLA models, the most robust one, RT-1-400k, succeeded in 45.6% of the test cases with the mutated camera poses that were passed with the default camera poses. OpenVLA-7b also passed 31.3% of the test cases. We noticed a significant performance gap between two Octo models and other VLA models, where both two Octo models only passed less than 10% of the test cases while the other five models all passed more than 20% of the test cases. These results may have been largely attributed to the fact that Octo models used the smaller training dataset (~ 65K robot demonstrations) compared with the other ones (130K ~ 160K). As a result, the Octo model’s generalizability to visual inputs captured from different camera angles is significantly affected.
Among seven subject VLA models, RT-1-400k was the most robust one against the perturbation of camera poses. Octo-series models performed significantly less robustly than the other models. This may be largely attributed to the fact that Octo models were trained with only half of the robot demonstration data compared with the other subject models.
5.5. RQ5: How Robust Do VLA Models Perform Against Unseen Objects?
To investigate this research question, we used an external object dataset, YCB (Calli et al., 2015). YCB contains 56 objects that are not included in the Open-Embodiment-X dataset (the training/fine-tuning dataset of our subject VLA models). We leveraged VLATest to generate 1,000 testing scenes for each of the four subject robotic manipulation tasks while sampling the target object and the obstacles from the YCB dataset. We compared the performance of these subject VLA models with objects from YCB dataset to the results in RQ1.
Table 5. RQ5) Performance of subject VLA models with seen/unseen objects. Each cell in the column Seen and Unseen represents the success rate of different VLA models in different tasks. Darker the color, larger the negative differences between the performance on seen and unseen objects.
VLA Models Task 1: Pick Up Task 2: Move Near Task 3: Put On Task 4: Put In
Seen Unseen Diff. Seen Unseen Diff. Seen Unseen Diff. Seen Unseen Diff.
RT-1-1k 0.7% 1.4% +100.0% 1.4% 0.9% -35.7% 0.0% 0.0% 0.0% 0.0%
RT-1-58k 28.2% 5.0% -82.3% 10.9% 3.9% -64.2% 0.0% 0.0% 0.0% 0.0%
RT-1-400k 34.4% 10.3% -70.1% 9.4% 2.9% -69.2% 0.5% 0.4% -20.0% 0.1% 0.1%
RT-1-X 19.5% 3.0% -84.6% 5.8% 1.3% -77.6% 2.3% 0.6% -73.9% 0.4% 0.7% +75.0%
Octo-small 0.8% 0.1% -88.9% 0.8% 1.5% +87.5% 2.2% 0.9% -59.1% 1.1% 0.8% -27.3%
Octo-base 0.0% 0.0% 0.6% 0.4% -33.3% 1.2% 0.2% -83.3% 1.1% 0.5% -54.5%
OpenVLA-7b 5.9% 3.5% -40.7% 12.7% 2.8% -78.0% 2.1% 0.9% -57.1% 1.1% 1.0% -0.9%
Avg. 12.8% 3.3% -74.2% 6.0% 2.0% -66.7% 1.2% 0.4% -66.7% 0.5% 0.4% -20.0%
  • $\dagger$ Diff. = (Unseen - Seen) / Seen
Table 5 shows the performance of seven subject VLA models when manipulating seen/unseen objects. Overall, when manipulating with unseen objects, the performance of VLA models dropped significantly compared with manipulating with seen objects. The average performance across different VLA models dropped by 74.2%, 66.7%, 66.7%, and 20.0% in Task 1, Task 2, Task 3, and Task 4, respectively. Notably, we also found that none of these VLA models completely failed when manipulating unseen objects (except those that also failed with seen objects). These results suggest that, while all VLA models exhibit certain possibilities of generalizing themselves to unseen objects, the performance is by far unreliable.
Current VLA models, though exhibit potential for generalization, are still unreliable for handling manipulation tasks with unseen objects. Our subject VLA models saw a performance drop ranging from 20.0% to 74.2% across four tasks when manipulating unseen objects compared with the seen objects.
We also examined each testing scene and the success of the VLA model at every individual step, similar to the investigation of RQ1 and RQ2, to understand how unseen objects might have impacted its performance. Specifically, since the model must succeed at each step sequentially to complete the testing scene, we define a metric called the transfer rate ($Tr$), which calculates the model’s ability to transfer success from one step to the next, as follows.
(1) $\footnotesize Tr_{n}=\frac{Success\_rate_{n}}{Success\_rate_{n-1}},$
where $Success\_rate_{n}$ denotes the success rate of a task at step $n$. We define $Success\_rate_{0}=1000$.
Figure 7. (RQ5) Transfer rate (TrTr) of each task when manipulating with seen/unseen objects. Entries with * mean that their mean differences are statistically significant ($p<0.05$).
We present the results in Fig. 7. We also performed a paired $t$-test to examine the statistical differences between $Tr$ of seen and unseen objects. Overall, though $Tr_{1}$, $Tr_{2}$, and $Tr_{3}$ all decreased when manipulating with unseen objects instead of the seen ones, we only saw significant statistical differences on $Tr_{1}$ in Task 1 and Task 2 ($p=0.011$ and $p=0.007$, Cohen’s $d$ effect sizes are $1.34$ and $0.891$, respectively). These results reveal that the biggest challenge for current VLA models when dealing with unseen objects remains to accurately locate the correct object to manipulate, suggesting that the VLA models may not be able to recognize unseen objects in many cases. We did not observe significant statistical differences on $Tr_{1}$ in Task 3 and Task 4. A plausible explanation is due to the small number of passed test cases (~ 20 out of 1000) in both tasks.
Current VLA models struggled to recognize unseen objects, which was the primary cause of failure in most test scenes.
6. Discussion
Our study reveals several key implications that can lead to the development of better and more reliable VLA models for robotic manipulation. In this section, we discuss these implications and explore future research opportunities.
VLA models for robotic manipulation—too good to be true for now. While VLA models have the potential to revolutionize AI-enabled robotic manipulation, our experiment results in RQ1 and RQ2 reveal that the current models are still unreliable for common robotic manipulation tasks at the time of this study. As a result, deploying VLA models in high-stakes and safety-critical applications remains impractical.
Our detailed analysis suggests that the deficiency mainly comes from the lack of capabilities in precisely interpreting complex task requirements and accurately localizing the correct target objects to manipulate. We believe there is a huge space for improvement in addressing these issues. One potential solution is to scale up the model size. Currently, even the largest model in our study, OpenVLA-7b, has only 7 billion parameters, which is significantly smaller than state-of-the-art (SOTA) models in other domains. For instance, one of the largest open-source LLMs, Llama 3.1, has 405 billion parameters. The closed-source LLMs such as GPT-4 and Claude-3.5 typically have even much higher numbers of model parameters. By scaling up the model, we may also observe emergent capabilities for VLA models (Brohan et al., 2023).
In addition to scaling up the size of VLA models, exploring more effective prompting strategies could be another promising direction. Previous studies have demonstrated that improved prompting can significantly enhance the performance of LLMs across various tasks, such as solving mathematical problems (Wei et al., 2022; Yao et al., 2024) and code generation (Zhang et al., 2023; Chen et al., 2023). Our Finding 5.1 suggests that current VLA models struggle to decompose complex task instructions into multiple steps of action. To fix this, one may consider prompting the VLA model step by step. This also aligns with one of the popular prompting strategies in the field of LLMs – chain-of-thought prompting. While our study has only explored the most basic prompts, future research should explore more advanced prompting techniques to evaluate their impact on VLA model performance. Furthermore, one can also introduce multi-agent systems to split the robotic manipulation tasks among multiple VLA agents, which is a strategy widely used in other domains (Xia et al., 2024; Zhang et al., 2024; Yang et al., 2024).
Addressing the robustness challenges. Our study results from RQ3 and RQ4 reveal that current VLA models lack robustness against several external factors, such as lighting conditions and camera poses. We also found that models with large-scale pre-training or those trained with larger robot demonstration datasets exhibit greater robustness compared to others. This highlights a promising research direction: enriching the robot demonstration data. Since manually collecting real-world robot demonstration data requires significant labeling efforts, researchers may leverage data augmentation techniques (Alomar et al., 2023; Bharadhwaj et al., 2024) or employing sim2real translation (Zhao et al., 2020; Höfer et al., 2021) to scale up training data by utilizing simulation environments.
Moreover, our Finding 5.3, Finding 5.3, and Finding 5.4 show that when external factors like lighting and camera angles deviate significantly from default settings, VLA model performance drops accordingly. This suggests that existing robot demonstration datasets may lack diversity, especially regarding variations in external conditions. In future work, when collecting datasets for training/fine-tuning factors, practitioners should take these environmental factors into account, which may lead to the training of more robust VLA models.
Assessing the capabilities of VLA models. Our Finding 5.5 and Finding 5.5 (RQ5) suggest that current VLA models struggled to perform tasks with unseen objects. In practice, it is unrealistic to expect VLA models to successfully perform manipulation for every possible task scenario. Thus, in parallel to the development of more powerful and robust VLA models, it is also important to design novel techniques to comprehensively assess the capabilities of VLA models and derive proper guidelines about the use of VLA models. To address these, potential solutions include throughout offline benchmarking and online risk assessment.
In terms of offline benchmarking, our 18,604 generated test scenes across four tasks can serve as one of the early benchmarks for VLA models. In future work, practitioners may focus on expanding this benchmark to include a broader range of robotic manipulation tasks and diverse robot settings, covering various object types, environmental factors, and task complexities. In terms of online risk assessment, one may consider adopting SOTA techniques for risk assessment in other domains, such as uncertainty estimation (Huang et al., 2023b; Mukherjee and Awadallah, 2020) and safety monitoring (Xie et al., 2024). While general techniques may be directly used in the context of robotic manipulation with VLA models, it is unclear to what extent they can help with assessing the quality and reliability of decisions made by VLA models.
Towards efficient testing for VLA models. Although our proposed VLATest successfully identified numerous failed test scenes across our subject VLA models, it also incurred significant time overheads. Specifically, we relied on a random sampler for sampling target objects and obstacles in the testing scenes, which may not have been the most efficient approach. Future research could focus on optimizing the $pose_{s}ampler$ in Algorithm 1 to strategically assign critical positions and orientations for objects and obstacles. For instance, one can consider metamorphic-based methods (Wang and Su, 2020; Yu et al., 2022; Gao et al., 2024) or search-based methods (Zhou et al., 2023; Cheng et al., 2023) to efficiently generate test scenes for VLA models. Meanwhile, researchers may also work on test prioritization (Li et al., 2024a; Dang et al., 2024; Feng et al., 2020) or test selection (Huang et al., 2024; Hu et al., 2023; Gao et al., 2022; Aghababaeyan et al., 2024) towards efficient testing for VLA models.
Nevertheless, as VLA models continue to evolve, we argue that testing strategies must also evolve in parallel. The complexity and capabilities of these models are likely to increase, requiring more sophisticated and adaptive testing frameworks that can keep pace with advancements.
7. Related Work
Foundation models for robotics. A large body of research has been done on the use of foundation models for robotics. Based on the data modality, recent work can be roughly categorized into two groups: (1) those that use large language models (LLMs) for robotics, and (2) those that use multi-modality foundation models (e.g., vision language models) for robotics. One of the early attempts in using LLMs for robotics is to leverage LLMs as a reward designer for Deep Reinforcement Learning-based robotic manipulation (Song et al., 2023b). This work proposes a self-refined framework to let the LLM automatically and interactively generate and refine reward functions used to train DRL policies for robotic control. Alternatively, Zhou et al. (Zhou et al., 2024) adopts LLMs, first, as a translator to convert natural language input into a Planning Domain Definition Language (PDDL) formulation and then outputs action plan sequences for hong-horizon tasks.
Different from the use of LLMs, multi-modality foundation models such as vision language models (VLMs) enable the possibility of advancing AI-enabled robotics in several novel tasks, e.g., for robotic manipulation (Brohan et al., 2022, 2023; Team et al., 2024), visual question answering with robots (Du et al., 2023), and visual state representations (Nair et al., 2022). Our work is most closely related to those using vision-language-action models, a specialized category of VLMs, for robotic manipulation (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2024c; Brohan et al., 2022; Li et al., 2023; Stone et al., 2023). One of the pioneer works in VLA models for robotic manipulation is RT-1 (Brohan et al., 2022), which uses a combination of a FiLM EfficientNet and a transformer to learn control policies from 130k real-world robot demonstrations. Since the publish of Open X-Embodiment dataset (Padalkar et al., 2023), a series of VLA models have been proposed by either training or fine-tuning on Open X-Embodiment (Brohan et al., 2023; Team et al., 2024; Kim et al., 2024; Li et al., 2024c). Our work is parallel to these works. We propose a general testing framework to test VLA models and present an empirical study to comprehensively assess the performance of VLA models.
Testing cyber-physical systems. Quality assurance of traditional cyber-physical systems (CPS) and AI-driven CPS is a vital topic in the realm of software engineering, which assures the safety and trustworthiness of such complex systems in safety-critical domains. Substantial efforts have been made by researchers and industrial practitioners to safeguard the quality of CPS from various perspectives, such as testing (Lee et al., 2023; Menghi et al., 2020; Chen et al., 2020; Ayerdi et al., 2021; Zhang et al., 2022), analysis (Stocco and Tonella, 2022; Biagiola and Tonella, 2024; Yohanandhan et al., 2020) and repairing (Song et al., 2023a; Wang et al., 2024; Guo et al., 2020). Considering robotic manipulation as a widely deployed representative CPS, our work is most related to the testing of CPS.
In particular, Lee et al. (Lee et al., 2023) propose MOTIF, a gray-box fuzzing-based approach to generate test data for software deployed in CPS. This method monitors the coverage achieved by the vanilla and the mutated functions, thereby exploring the behaviour space of the software under test. Differing from the existing testing solutions, MOTIF specifically targets the testing challenges of C and C++ languages that are widely used in CPS contexts. In terms of the black-box approach, Menghi et al. (Menghi et al., 2020) introduces ARIsTEO, a searched-based testing method with a loop of approximation and refinement to identify requirement violations of CPS models, and Chen et al. (Chen et al., 2020) propose an active fuzzing approach to find test suits for packet-level CPS network attacks. Moving forward with CPS with AI module embedded, Zhang et al. (Zhang et al., 2022) design a series of time-aware coverage criteria for DNN controllers in CPS and develop a falsification framework FalsifAI, to find system defects by leveraging the coverage information from the proposed criteria. It is worth noting that the aforementioned works are not well applicable to VLA models, which is attributable to the multimodality characteristic, autoregressive generation mechanism, and large-scale model size. In contrast, our study initiates a testing framework tailored for VLA models in the context of different robotic manipulation tasks. VLATest encompasses four multimodality-aware testing operators to assess the VLA models’ performance from different perspectives.
Benchmarking and testing foundation models. After the remarkable success of foundation models, an important research direction is to explore and understand the boundaries of their capabilities, which can provide a basis for further safeguarding and enhancement. Based on this motivation, numerous studies have focused on evaluating large language models in the text domain across various properties, such as correctness (Liang et al., 2023a), factuality (Lin et al., 2022; Dziri et al., 2022), robustness (Zhu et al., 2023; Wang et al., 2023), fairness (Sun et al., 2024), and privacy (Hamid et al., 2023). These studies offer a wealth of resources for the general performance assessment of LLMs. On the other hand, researchers have also made significant efforts to evaluate specific capabilities of LLMs, such as code-related abilities. These efforts include validating the correctness of generated code (Cassano et al., 2023; Liu et al., 2024b; Yu et al., 2024), examining whether LLMs can address real-world GitHub issues (Jimenez et al., 2023), and analyzing error patterns in LLM-driven code translations (Pan et al., 2024).
While the aforementioned studies provide valuable insights, most of them focus on performing analysis on static benchmarks. Orthogonal to these related works, there are also several attempts at dynamically generating test cases to evaluate one or more key properties of models. For example, existing studies use rule-based methods to generate test cases that measure bias (Wan et al., 2023) or linguistic capabilities (Lee et al., 2024), employ metamorphic testing for evaluating language translation (Sun et al., 2020; He et al., 2020), and utilize mutation-based frameworks to assess robustness (Xiao et al., 2023). Our work differs from these previous approaches in two main ways: (1) our framework centers on VLA, a novel architecture that integrates multi-modal inputs and generates complex control commands, distinguishing it from language-focused foundation models; (2) our framework generates test cases that account for complex interactions in robotic manipulation within a 3D environment, presenting a new challenge compared to prior text-only testing frameworks.
8. Threats to Validity
We discuss the threats to the validity of our empirical study results, as well as the mitigating factors.
Internal Validity. One potential threat comes from the randomness in our experiments. To mitigate this, we generated 18,604 testing scenes in our empirical study. Overall, our empirical study took over 400 GPU hours to finish. While we only executed each testing scene once per VLA model, we believe this has limited impact on our findings as our goal is to evaluate the VLA model’s overall performance on each task instead of the success of one specific testing scene.
Another potential threat to internal validity is the choice of prompt templates. To mitigate this, we followed the previous works (Team et al., 2024; Kim et al., 2024; Brohan et al., 2022, 2023) to use the standard prompt template when evaluating VLA models for robotic manipulation. We also discuss the impact of prompts and the potential of improving VLA models with better prompting in Sec. 6.
External Validity. Threats to external validity include whether our study results can generalize to different experimental settings, such as different robotic manipulation tasks and different VLA models. To mitigate these threats, we select the four most popular robotic manipulation tasks according to Open X-Embodiment dataset (Padalkar et al., 2023). In terms of the VLA models, we included the SOTA publicly available VLA models by the time we conducted this study. Given that the VLA model for robotic manipulation is a fast-developing research area, we also plan to extend our evaluation to other SOTA VLA models, e.g., RT-2 (Brohan et al., 2023), once they are made available.
Finally, we have only implemented VLATest and evaluated the VLA models within one simulation environment, Maniskill2 (Gu et al., 2023). To reduce the impact of distribution shifts between simulation and real robot execution, we chose a improved version of Maniskill2 (Li et al., 2024b), which provides visual matching to render realistic images as the visual input for VLA models.
Construct Validity. When designing our testing framework, VLATest, we have only considered a limited number of testing operators. To combat the threat, we included testing operators covering different aspects, e.g., task difficulties (i.e., number of obstacles) and environmental factors (i.e., the lighting intensity and camera poses). In future work, one may continue to improve VLATest by adding more testing operators, such as the number of lighting sources, the intrinsic parameters of cameras, and the resolution of cameras, for more comprehensive evaluation of VLA models.
9. Conclusion
In this paper, we propose VLATest, one of the early testing frameworks for testing VLA models for robotic manipulation. VLATest is a generation-based fuzzing framework based on ten operators covering different perspectives in the test scene of robotic manipulation. Upon implementing VLATest with Maniskill2 simulation environments, we further conducted a large-scale empirical study to assess the performance and robustness of seven popular VLA models across four robotic manipulation tasks. We generated 18,604 testing scenes, conducted more than 400 GPU hours of simulation and performed a detailed analysis of the challenges and limitations faced by the current VLA models. At the end of the paper, we discuss the implications of our study, shedding light on several future research directions to improve the quality and reliability of VLA models for robotic manipulation.